### v1

# WINTER on S250206dm: A near-infrared search for an electromagnetic counterpart

Abstract
We present near-infrared follow-up observations of the International Gravitational Wave Network (IGWN) event S250206dm with the Wide-Field Infrared Transient Explorer (WINTER). WINTER is a near-infrared time-domain survey designed for electromagnetic follow-up of gravitational-wave sources localized to 300 deg2. The instrument’s wide field of view (1.2 deg2), dedicated 1-m robotic telescope, and near-infrared coverage (0.9-1.7 microns) are optimized for searching for kilonovae, which are expected to exhibit a relatively long-lived near-infrared component. S250206dm is the only neutron star merger in the fourth observing run (to date) localized to 300,deg2 with a False Alarm Rate below one per year. It has a probability of being a neutron star–black hole (NSBH) merger and a probability of being a binary neutron star (BNS) merger, with a credible region spanning 38 deg2, an estimated distance of 373 Mpc, and an overall false alarm rate of approximately one in 25 years. WINTER covered of the probability area at least once and at least three times. Through automated and human candidate vetting, all transient candidates found in WINTER coverage were rejected as kilonova candidates. Unsurprisingly, given the large estimated distance of 373 Mpc, the WINTER upper limits do not constrain kilonova models. This study highlights the promise of systematic infrared searches and the need for future wider and deeper infrared surveys.
1 Introduction
The discovery of GW170817 during the second Advanced LIGO-Virgo observing run (O2) marked a breakthrough in multi-messenger astronomy, providing the first confirmed detection of a kilonova powered by r-process nucleosynthesis in the neutron-rich ejecta of a binary neutron star (BNS) merger (Abbott et al., 2017a, c, b). This event established the importance of joint observations of gravitational wave (GW) alerts and identifying electromagnetic (EM) counterparts and investigating the physics of compact object mergers, the neutron star equation of state, and heavy-element production (e.g., Abbott et al. 2017b; Barnes et al. 2016; Coulter et al. 2017; Barnes & Kasen 2013; Evans et al. 2017; Goldstein et al. 2017; Grossman et al. 2014; Haggard et al. 2017; Hallinan et al. 2017; Kasen et al. 2017, 2013; Kasliwal et al. 2019; Li & Paczynski 1998; Margutti et al. 2017; Metzger et al. 2010; Metzger 2020; Roberts et al. 2011; Rosswog 2005; Tanaka & Hotokezaka 2013; Tanvir et al. 2017; Troja et al. 2017).
Despite extensive optical and infrared follow-up campaigns, subsequent LIGO-Virgo observing runs (e.g., the third observing run, O3) yielded no confirmed EM counterparts to BNS or neutron star-black hole (NSBH) mergers (e.g., Coughlin et al., 2020; Kasliwal et al., 2020; Abbott et al., 2021a, b; Collaboration et al., 2021; Zhu et al., 2021a). Many of these events were predicted to be too faint for detection by optical telescopes, as kilonovae exhibit viewing-angle-dependent blue emission that fades rapidly (1 week), while a longer-lived near-infrared (NIR) signal is predicted to be more isotropic (Kasen et al., 2013, 2017; Metzger, 2020). Indeed, GW170817 had significantly longer-lived infrared emission compared to optical emission (Tanvir et al., 2017). Zhu et al. (2021b) predicts kilonovae may be up to 8–10 times more detectable in the near-infrared compared to optical wavelengths. However, systematic sensitive wide-field infrared searches have been limited by high detector costs and high sky backgrounds. Some shallow searches were performed by the Palomar Gattini-IR survey (Moore et al., 2016; De et al., 2020) during O3 (Coughlin et al., 2019). Two new wide-field IR surveyors have come online in the fourth observing run, including the Wide-field Infrared Transient Explorer (WINTER), discussed in this study, and PRIME a new NIR survey on a 1.8-meter telescope with four large-format HgCdTe sensors covering a 1.56 deg2 field of view (Durbak et al., 2024; Kutyrev et al., 2023; Yama et al., 2023).
WINTER is a 1-meter telescope and camera at Palomar Observatory designed for rapid-response NIR transient discovery, with a wide field of view () and infrared sensitivity ( mag) (Lourie et al., 2020). WINTER’s InGaAs-based detectors provide a cost-effective alternative to traditional HgCdTe NIR sensors, enabling systematic follow-ua of GW events in , , and shortened- bands (0.9-1.7 microns) (Simcoe et al., 2019).
WINTER is designed around the discovery of EM counterparts to GW events, with a wide field of view, infrared sensitivity, and rapid robotic follow-up to enable efficient tiling of GW localization regions (Frostig et al., 2020). Frostig et al. (2022a) presents a comprehensive simulation of WINTER’s capability to follow up BNS mergers during the fourth International Gravitational Wave Network (IGWN) observing run (O4). The analysis predicts that near-infrared kilonovae will persist longer than their optical counterparts, with red kilonovae detectable at distances approximately 1.5 times greater in the NIR. This study offers detailed projections of WINTER’s effectiveness in identifying new kilonovae. However, issues manufacturing the new-to-astronomy InGaAs detectors resulted in a decreased final sensitivity ( instead of magnitudes), impacting WINTER’s effectiveness following up kilonovae (Frostig et al., 2024).
In this paper, we present results from WINTER follow-up of the LIGO–Virgo–KAGRA (LVK; Aasi et al. 2015; Acernese et al. 2014; Somiya 2012; Aso et al. 2013; Akutsu et al. 2020) gravitational wave trigger S250206dm (LIGO–Virgo–KAGRA Collaboration, 2025a). S250206dm was first detected by the LVK at 2025-02-06T21:25:30.439 UT, as part of the ongoing O4 run. The event was recovered by the GW search pipelines GstLAL (Messick et al., 2017; Sachdev et al., 2019; Hanna et al., 2020; Cannon et al., 2021; Ewing et al., 2024; Tsukada et al., 2023), MBTA (Adams et al., 2016; Aubin et al., 2021) and PyCBC Live (Allen et al., 2012; Allen, 2005; Dal Canton et al., 2020; Usman et al., 2016; Nitz et al., 2017; Davies et al., 2020), with an estimated distance of 373 Mpc and a false alarm rate of about 1 in 25 years. The event, as reported by the LVK, has a 55% probability of being a NSBH merger, with a 37% probability of being a BNS merger (LIGO–Virgo–KAGRA Collaboration, 2025a); in addition, the probability that the lighter compact object is a neutron star is (LIGO–Virgo–KAGRA Collaboration, 2025b). S250206dm’s localization underwent several revisions, with the latest (released 36 hours after merger) having a 50% localization area of 38 deg2, and a 90% area of 547 deg2. With its relatively precise localization and confident detection, S250206dm is one of the most promising merger candidates in O4 with a high probability of containing a neutron star.
In addition to the gravitational wave properties of S250206dm which suggested a possible electromagnetic counterpart, the IceCube Neutrino Observatory (Aartsen et al., 2017) promptly reported the detection of two neutrinos in spatial and temporal coincidence with the event (IceCube Collaboration, 2025), as part of its realtime search for neutrino-gravitational wave coincident detections (see e.g Abbasi et al., 2023). Neutrino 1 was detected 109s before merger, while neutrino B was detected 289s after merger.111https://roc.icecube.wisc.edu/public/LvkNuTrackSearch/output/2025_02_06_S250206dm-7-Update.html We performed additional ToO observations of neutrino fields, similar to other teams (Becerra et al., 2025; Masi, 2025).
In this paper, we report on the nondetection of an EM counterpart to S250206dm, observed with WINTER during O4. We describe our observational strategy, data processing, and implications for future NIR follow-up of GW events.
2 WINTER follow-up campaign
2.1 WINTER observatory
The WINTER camera is mounted on a dedicated 1 m telescope at Palomar Observatory, and comprises six InGaAs detectors that together produce a field-of-view of 1.2 deg. x 1 deg with a pixel scale of 1.13″per pixel. The WINTER telescope has been running robotically on sky since June 2023 (Frostig et al., 2024), featuring new InGaAs sensor technology and a custom read out (Frostig et al., 2022b; Malonis et al., 2020), a novel fly’s-eye optical design (Lourie et al., 2020), and custom opto-mechanics (Hinrichsen et al., 2020).
The WINTER telescope is operated by the WINTER Supervisory Program (WSP), a multithreaded python-based control system that manages scheduling and observations using PyQt5 for interfacing with the Qt framework. Daemon processes on the observatory’s control PC initiate WSP, which automates key observatory functions, including nightly scheduling, target-of-opportunity (ToO) handling, and priority-based command execution, while allowing manual overrides when necessary. Drawing on methodologies from MINERVA (Swift et al., 2015), Robo-AO (Riddle et al., 2012), and GOTO (Dyer et al., 2018), WSP facilitates automated decision-making for telescope operations, including dome control and observation execution. Target of Opportunity (ToO) observations are submitted automatically through a dedicated API system, which validates scheduling requests and sends them to the scheduler.
2.2 S250206dm observations
WINTER observations of the GW localization region commenced on UT 2025-02-08T04:05:52, 30.5 hours after the merger. No observations were possible on the first night after the merger due to poor weather at Palomar Observatory, which continued to limit observations throughout the follow-up campaign. WINTER observations are conducted on a predefined field grid. An optimal schedule to tile the GW skymap using WINTER fields was computed using gwemopt (Coughlin et al., 2018, 2019; Dietrich et al., 2020). Each WINTER field was observed in the J-band with a set of eight two-minute long dithered exposures, corresponding to a combined integration time of sixteen minutes per field. These fields were subsequently reobserved on the nights of UT 2025 Feb 9, 10, 11, 16, and 17. All WINTER pointings overlaid on the GW-skymap are shown in Figure 1. The full details of the pointings are available on TreasureMap (Wyatt et al., 2020). Within ten days of merger, a total of 43% of the probability enclosed within the latest available GW skymap at the time of writing222https://gracedb.ligo.org/api/superevents/S250206dm/files/Bilby.offline1.multiorder.fits was observed with WINTER at least once, 40% of the total probability was observed at least twice, and 35% was observed at least three times.
2.3 Data processing and transient detection
The WINTER images were processed by the WINTER data reduction pipeline implemented in mirar333https://github.com/winter-telescope/mirar (Karambelkar & Stein et al., in prep) – a modular, python-based framework designed for real-time, end-to-end processing of astronomical images to search for transients. A detailed description of the processing steps will be provided in a forthcoming publication. Briefly, the pipeline begins by performing dark subtraction, flat calibration and sky-subtraction on the raw WINTER images. Astrometric solutions are then computed using the packages Astrometry.net (Lang et al., 2010) and Scamp (Bertin, 2006). Eight astrometry-calibrated dithers are then stacked using the software Swarp (Bertin et al., 2002), and the stacked images are photometrically calibrated relative to the Two Micron All-Sky Survey (2MASS; Skrutskie et al., 2006) point source catalog (Cutri et al., 2003). Images from each of WINTER’s six detectors are reduced separately, so each sixteen-minute observation of a field produces six stacked images each with dimensions 0.6 deg x 0.3 deg. The median depths achieved during our observations of S250602dm for the six detectors were 17.6, 17.6, 17.4, 18.0, 17.5, and 17.4 mag (AB) respectively. Of the total GW skymap probability, 41% was observed to a depth of 17 mag or deeper, 38% was observed to a depth of 17.5 mag or deeper, 31% was observed to a depth of 18 mag or deeper, and 11% was observed to a depth of 18.5 mag or deeper.
Image subtraction is then performed on these stacked images relative to archival J-band reference images from the United Kingdom InfraRed Telescope (UKIRT) Hemisphere Survey (UHS; Dye et al., 2018) using the optimal image-subtraction algorithm ZOGY (Zackay et al., 2016). Sources identified in the difference image are cataloged in a SQL database, cross-matched to external point-source catalogs such as 2MASS, Pan-STARRS (PS1; Chambers et al., 2016), as well as to transient alerts from ZTF and the Transient Name Server, packaged as alerts in the avro format, and broadcast to the Fritz.science instance of Skyportal for scanning and visual vetting (Coughlin et al., 2023).
During the ten day observational campaign, the WINTER images were processed every night in real-time and the alerts were examined to identify interesting transients. Separately, at the end of the campaign, we reprocessed the WINTER images by stacking repeat observations of the same field from multiple nights to increase the depth of our search. The median depths for these “super-stacks" was 17.8, 17.7, 17.7, 18.0, 17.8, and 17.5 mag for the six detectors respectively.
2.4 Candidate Searches
A total of 62991 individual alerts were generated by our standard observations, of which 62675 lay within the 95% contour reported by LVK. We perform three independent and complementary searches for possible transients in our data and reduce the number of false positive results.
2.4.1 Candidates With Multiple WINTER detections
Our data reduction pipeline cross-matches each new detection to a dynamically-updated table of WINTER ‘sources’, with each source having an assigned name. The new detections are then crossmatched to all previous detections of the same source, with each avro alert including a complete history of detections in our data. The 62675 alerts were grouped into 57670 unique sources. We select those sources with multiple detections, in order to reject moving objects and reduce the number of bogus sources. Of 57670 sources, we select the 4141 sources with at least 2 detections.
2.4.2 Candidates with a ZTF crossmatch
Every WINTER detection is crossmatched to all archival ZTF alerts (Bellm et al., 2018), providing a ZTF source name if available. 9608 of the WINTER sources have a pre-assigned ZTF name. For this event, we perform an additional crossmatch on all ZTF alerts detected in the same 2 week period post-merger, providing additional matches in cases where the WINTER detection precede ZTF detections. A total of 116161 ZTF alerts were detected within 14 days of merger in the 95% contour, for which no additional cuts were applied. These including both ToO and serendiptious detections of the ZTF public survey. The additional crossmatch yields 1275 matches, of which 65 are new. In total, 9673/57670 WINTER sources have a ZTF crossmatch.
Name | RA | Dec | Date | Limiting |
Magnitude [AB] | ||||
AT2025bmq | 38.518 | 54.572 | 2025-02-08T06:44:50.635 | 18.3 |
2025-02-10T05:15:34.858 | 17.7 | |||
AT2025bbp | 153.782 | -22.882 | 2025-02-08T11:04:14.707 | 17.3 |
2025-02-11T09:29:20.609 | 17.0 | |||
AT2025ban | 152.918 | -18.952 | 2025-02-09T08:58:45.274 | 14.4 |
AT2025baj | 156.517 | -28.346 | 2025-02-08T11:21:47.004 | 14.6 |
2025-02-11T10:39:07.412 | 17.3 |
2.4.3 Candidates with a TNS cross-match
We select all candidates included on the TNS event page for S250206dm, which was 116 at the time of writing.444https://www.wis-tns.org/system/files/ligo/o4/S250206dm_20250206_212530/S250206dm_20250206_212530_AFTER.json. Of these, 95 lie within the 95% contour, though just four TNS transients fell within fields observed by WINTER. We crossmatch all WINTER candidates to the TNS transients with a radius of 3", but find no coincident detections.
In Table 1 we list each WINTER image which overlapped with a TNS transient. Of the four TNS transients, only AT2025bmq fell in a region observed by WINTER and had -band reference images, both necessary requirements to generate an alert. As noted in the table, no such alert was produced. AT2025bmq is coincident with an IR-bright host which is clearly recovered in the WINTER images at . This is consistent with the archival magnitude of this source, so we cannot resolve whether there is any excess flux from a nuclear transient.
For the other three sources (AT2025bbp, AT2025ban and AT2025baj), no reference images were available, so no image subtraction was performed. All three have apparent underlying hosts, but given the depths of our images, neither the host nor transient were recovered in our science images. Only one of these three transients (AT2025baj) has a host bright enough to be detected in 2MASS. However, at =17.4 mag, this host is fainter than the limiting magnitudes of our science images.
WINTER + WINTER | WINTER + ZTF | WINTER + TNS | Combined Total | |
Initial Sources | 4141 | 9673 | 0 | 11802 |
After Stellar Cuts | 140 | 140 | 0 | 274 |
After WINTER predection cuts | 140 | 140 | 0 | 274 |
After ZTF predection cuts | 135 | 4 | 0 | 139 |
After human vetting | 0 | 0 | 0 | 0 |
2.4.4 Candidate Vetting
Though many sources remain at this stage, the vast majority will be stellar. For all 11802 candidates selected by our three independent methods, we perform a series of algorithmic cuts:
We crossmatch all sources to PS1 to a star/galaxy morphology classifier trained using PS1 data (Tachibana & Miller, 2018). We reject WINTER sources within 7" of bright (m) stars, and those within 3" of fainter stars.
We reject sources within 3" of sources detected in Gaia-DR3 (Gaia Collaboration et al., 2023) with parallax.
We finally remove sources within 20" of very bright Gaia-DR3 stars, defined as M.
We reject sources which are ‘old’, defined as having either a WINTER or ZTF detection prior to merger.
For sources passing these cuts, we perform visual inspection. We were aided by crossmatches to Gaia-DR3 (Gaia Collaboration et al., 2023), SDSS (York et al., 2000) and WISE (Wright et al., 2010) for identifying known variable sources. We reject candidates based on the following criteria:
Bogus sources with poor subtraction residuals, especially those detections close to poorly-subtracted reference image sources. We require candidates to have at least two detections which pass visual inspection and appear to be plausibly real.
On this basis, every WINTER candidate is rejected as a possible transient. A full summary of cut statistics is given in Table 2.
3 Comparison with kilonova models
We generate model light curves for kilonovae using NSBH and BNS models presented in Ahumada et al., in prep. These models are simulated with the latest version (Bulla, 2023) of the 3D Monte Carlo radiative transfer code possis (Bulla, 2019) assuming a two-component (dynamical+wind) ejecta configuration and using state-of-the-art heating rates (Rosswog & Korobkin, 2024) and opacities (Tanaka et al., 2020). The kilonova grids span the expected parameter space, with the NSBH grid varying the average velocity of the dynamical ejecta and the masses of the dynamical and wind ejecta (for a total of 108 models) and the BNS grid the mass, average velocity and average electron fraction for both components (for a total of 3072 models). To account for distance uncertainties, we sample 100 distances from the probability distribution function for S250206dm, as provided by the multi-order Bilby skymap (LIGO–Virgo–KAGRA Collaboration, 2025b). For each sampled distance, we compute a full model grid of calculated ejecta properties to explore the range of possible emission scenarios.
Figure 2 presents the resulting marginalized model prediction, obtained by interpolating all light curves onto a common time grid and computing summary statistics at each epoch. We present the -band median light curve and confidence intervals and compare to the -band median light curve to demonstrate the relatively bright and persistent NIR plateau. Additionally, we overlay the limiting magnitudes of all WINTER follow-up observations (as described in Section 2) to assess whether the models are observationally constrained.
The absence of a detected kilonova in WINTER observations is consistent with the predicted NSBH and BNS model grids. At the estimated distance of S250206dm and the timing of the observations, the WINTER observations are not sufficiently deep to place meaningful constraints on any of the kilonova models explored in this study.
The WINTER’s effectiveness in gravitational-wave counterpart searches depends on both weather conditions and achieved sensitivity. The follow-up campaign for S250206dm was hindered by poor weather, which prevented early observations on the first night after the GW trigger. Under clear conditions, WINTER can respond robotically to a trigger in under ten minutes. Had WINTER been able to observe S250206dm within the first hour, it could have constrained some of the NSBH models presented here within the 3 credible interval, though it would still lack the depth to constrain any BNS models or a typical NSBH merger at 373 Mpc. In clear observing conditions, WINTER can achieve stacked -band depths of 19.5 mag, compared to only 18.0 mag in this weather-limited campaign. This difference significantly impacts the search horizon: the peak of the median NSBH (BNS) model shown would be detectable out to 219 (74) Mpc, compared to just 109 (37) Mpc under poorer conditions. Furthermore, issues with manufacturing the WINTER sensors lead to a significant decrease in achievable depth from the design (see Frostig et al. (2024) for more details). At the designed limiting magnitude of = 21 mag, WINTER would be able to constrain the peak of the median NSBH (BNS) merger model to 437 (149) Mpc.
4 Summary and way forward
In this study, we present results from the WINTER’s NIR follow-up observations of the gravitational wave event S250206dm, a candidate NSBH or BNS merger detected during the fourth LIGO-Virgo-KAGRA GW observing run. We conducted a targeted follow-up campaign to search for a potential EM counterpart. With systematic coverage of of the probability region within ten days post-merger, we recover no EM counterpart.
The nondetection is consistent with the modeled light curves of kilonovae and our observational depths (median mag by WINTER sensor), which were insufficient to place strong constraints on kilonova models at these distances. With more typical weather conditions and reduced latency—particularly if follow-up had occurred within 1–2 hours post-merger—WINTER could have been sensitive to the brighter end of predicted models, potentially placing meaningful constraints. Furthermore, no confirmed counterpart was reported by other optical or infrared facilities, highlighting the challenges of detecting EM counterparts to distant compact object mergers with current facilities.
This study underscores the need for deeper infrared follow-up capabilities to improve kilonova detection prospects. WINTER, as a dedicated NIR transient survey, provides an important step toward systematic NIR follow-up of gravitational wave events. However, the development of future instruments with greater sensitivity, wider fields of view, and improved sky coverage will be critical for detecting and characterizing kilonovae from NSBH and BNS mergers. The upcoming Cryoscope surveyor in Antarctica will cover a 50 field of view in the -dark to a depth of 24.1 mag AB in one hour (Kasliwal et al., 2025) and the Nancy Grace Roman Space Telescope will provide a depth of 27.60 mag in an hour in the F129 filter over a field of 0.281 .555https://roman.gsfc.nasa.gov/science/WFI_technical.html Furthermore, more precise and rapid gravitational wave sky localizations from future observing runs will improve targeting strategies for electromagnetic counterpart searches. Finally, next-generation ground-based observatories, such as the Vera C. Rubin Observatory (Ivezić et al., 2019), will offer complementary optical coverage, further expanding the capability to identify and study transient astrophysical phenomena.
Although WINTER did not detect a counterpart S250206dm, this follow-up campaign demonstrates a systematic NIR multi-messenger approach. Future attempts, coupled with advancing technology, will further improve our odds of identifying and characterizing the electromagnetic signatures of compact object mergers, deepening our understanding of neutron star physics, nucleosynthesis, and the transient universe.
Acknowledgments
WINTER’s construction is made possible by the National Science Foundation under MRI grant number AST-1828470 with early operations supported by AST-1828470. Significant support for WINTER also comes from the California Institute of Technology, the Caltech Optical Observatories, the Bruno Rossi Fund of the MIT Kavli Institute for Astrophysics and Space Research, the David and Lucille Packard Foundation, and the MIT Department of Physics and School of Science. D.F.’s contribution to this material is based upon work supported by the National Science Foundation under Award No. AST-2401779. This research award is partially funded by a generous gift of Charles Simonyi to the NSF Division of Astronomical Sciences. The award is made in recognition of significant contributions to Rubin Observatory’s Legacy Survey of Space and Time. M.B. acknowledges the Department of Physics and Earth Science of the University of Ferrara for the financial support through the FIRD 2024 grant.